# Remasking Discrete Diffusion Models with Inference-Time Scaling

### v1

@@ 6 Discussion and Related Works
Denoising Diffusion Implicit Models  An analogy can be drawn between our work and DDIM (Song et al., 2020a), where ReMDM is to MDLM for discrete signals as DDIM is to DDPM (Ho et al., 2020) for continuous data. That is, stemming from similar insights about the NELBO depending only on marginals, both our work and DDIM propose alternative forward and backward processes that maintain marginals while enabling more flexible generation relative to previous state-of-the-art diffusion models. Of note, Song et al. (2020a) also present a way of adapting non-Markovian processes to discrete domains; however, they focus on uniform categorical as the limiting distribution. In Appendix B, we demonstrate how this can be adapted for absorbing state diffusion and derive an equivalence to our method.
@@
Discrete Predictor-Corrector Samplers  Continuous time Markov chain (CTMC) theory provides a framework for samplers that correct errors in the reverse process, extending the original predictor-corrector formulation for continuous data (Song et al., 2020b) to discrete domains. In addition to the FB and DFM correctors discussed above, other correctors include: the Stein operator (Zhao et al., 2024), and DPC (Lezama et al., 2023) which improves the sample quality of MaskGiT by corrector sampling. Unlike the plug-and-play methods of FB, DFM, and ReMDM, the Stein operator and DPC correctors require additional training.
Although ReMDM sampler can be reformulated as equivalent to the DFM corrector sampler (Proposition 4.3), ReMDM outperforms this method (Table 1). For DFM, we use the hyperparameters that were found to produce best results in Gat et al. (2024). However, unlike ReMDM the best DFM schedule was not designed with a focus on remasking and when to utilize it. In Figure 5, we depict the remasking schedule corresponding to results in Table 1 ($T=4096$) for DFM and ReMDM, an illustrative example of this claim (see full results in Appendix E.2.1). The DFM schedule demonstrates a spike in the beginning and sharp decay afterward. In contrast, the ReMDM-loop schedule is designed to provide non-trivial remasking probability after a good ‘candidate’ sequence has been generated by standard MDLM generation. Additionally, ReMDM possesses other theoretical advantages over DFM, namely the bound on $\sigma_{t}$ and the NELBO analysis of the underlying probabilistic generative models that are not present in Gat et al. (2024).
Figure 5: ReMDM and DFM $\sigma_{t}$ schedules for OWT.
@@ 7 Conclusion
In this work, we have presented a novel family of absorbing state discrete diffusion samplers. Our method leverages the strong language modeling performance of this class of models by enabling the use of pretrained weights with the added benefit of more flexible sampling strategies that allow for remasking of predicted tokens. We demonstrate empirically that this leads to improved sample quality for both unconditional and conditional generation and leads to better controlled generation. Our approach also unlocks an important inference-time compute scaling axis that is more limited for existing masked diffusion models.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, specifically those related to the generation of synthetic text and images. Our work can also be applied to the design of biological sequences, which carries both potential benefits and risks.
@@ A.4 Deriving the ReMDM NELBO
The variational objective or negative evidence lower bound (NELBO) for diffusion models has the following form (Sohl-Dickstein et al., 2015):
@@
Starting from (A.4), we note that, in practice, we set $p_{\theta}(\mathbf{z}_{1})=q(\mathbf{z}_{1}\mid x)=q(\mathbf{z}_{1})=\bm{\pi}$ which ensures $\mathcal{L}_{prior}=0$. Additionally, the reconstruction loss $\mathcal{L}_{recon}$ is equivalent in both (A.4) and (11). We therefore turn our attention to the ReMDM diffusion loss term
@@
Additionally, we note that in Sahoo et al. (2024), MDLM is parameterized using the denoising $\mathbf{x}_{\theta}$ that is restricted in two ways. First the denoising model places zero probability mass on the [MASK] token, i.e., $\mathbf{x}_{\theta}^{\top}\bm{m}=0.$ Second the model ‘carries-over’ unmasked tokens, i.e., $\mathbf{x}_{\theta}^{\top}\mathbf{z}_{t}=1,$ if $\mathbf{z}_{t}\neq\bm{m}.$ Below we assume that our denoising network follows a similar parameterization.
@@
Then rewriting the summation in (24) as an expectation with $t$ sampled uniformly from $\{\frac{1}{T},\ldots,1\}$ and plugging in (27), we recover the diffusion loss term in (11).
@@
The forward-backward corrector sampler (Campbell et al., 2022) is derived using continuous time Markov chain theory. In particular, the rate matrix for the corrector sampler step $\mathbf{R}_{corrector}$ is the sum of the forward rate matrix $\mathbf{R}_{t}$ and the backward rate matrix $\tilde{\mathbf{R}}_{t}$, i.e. $\mathbf{R}_{corrector}=\mathbf{R}_{t}+\tilde{\mathbf{R}_{t}}$.
@@
From Sahoo et al. (2024), we know that the forward step of MDLM $q(\mathbf{z}_{t}\mid\mathbf{z}_{s})=\mathrm{Cat}(\mathbf{z}_{t};\frac{\alpha_{t}}{\alpha_{s}}\mathbf{z}_{s}+(1-\frac{\alpha_{t}}{\alpha_{s}})\bm{m})$. Let $\mathcal{V}$ denote the vocabulary set.
@@
The DFM corrector sampler (Gat et al., 2024) is derived from a generating velocity $u_{t}^{corr}$ using the following transformation equation.
@@
Note that our formulation is slightly different from the original presentation in Gat et al. (2024), since in our notation as $t$ goes from 1 to 0, we move from noise to the target distribution, whereas in the flow matching literature this direction is reversed. Plugging (49) and (50) into (48), we derive the following form for $u_{t}^{corr}$:
@@ Appendix B Comparison to DDIM Non-Markovian Processes
In DDIM (Song et al., 2020a), the authors present non-Markovian forward processes for both continuous and discrete signals. For discrete data, although in DDIM the proposed method assumes a uniform categorical as the limiting distribution, here we demonstrate how one can adapt the proposed method in DDIM for absorbing state diffusion and derive an equivalence between ReMDM and this new DDIM process under a reparameterization of $\sigma_{t}$. For clarity, we use $\sigma_{t}^{DDIM}$ to denote the parameter from DDIM and $\sigma_{t}^{ReMDM}$ to denote the parameter used in our work.
@@ Appendix C ReMDM Sampler Algorithms
Appendix D Additional Experimental Details
D.1 OpenWebText
In this experiment, we reuse the pretrained AR, SEDD, and MDLM checkpoints released by (Sahoo et al., 2024) where the diffusion models are trained using a log-linear schedule, i.e., $\alpha_{t}=1-t$. AR, SEDD, and MDLM share the same architecture: a Transformer-based model (Vaswani, 2017) that augments the diffusion transformer (Peebles & Xie, 2023) with rotary embeddings (Su et al., 2024) and consists of 169M parameters. The neural network is comprised of 12 layers, 12 attention heads, and 768 hidden dimensions. Please see Sahoo et al. (2024) for the full model architecture and training details.
As in Sahoo et al. (2024), we use gpt-2 tokenizer for our experiments. We use the same train-validation split as in Sahoo et al. (2024) (where the last 100k documents of OWT were designated as the validation set) and randomly select 5,000 samples from the validation set to serve as the ‘reference’ for MAUVE score computation. For the discrete flow matching corrector sampler, we follow Gat et al. (2024) and set the corrector schedule $\beta(t)=At^{0.25}(1-t)^{0.25}$ where $A=10$ (see Appendix A.7 for the definition of corrector schedule). We empirically find that nucleus sampling performs better than the temperature technique proposed in Gat et al. (2024). We also use non-fixed-width time steps as in Gat et al. (2024).
For evaluation metrics, we report MAUVE scores, generative perplexity, and entropy. For MAUVE, we generate 5,000 samples for each model/sampler. We use the gpt-2 tokenizer, GPT-2 Large (Radford et al., 2019) as the embedding model, and the MAUVE scaling hyperparameter is set to 5. For generative perplexity, we use GPT-2 Large as the external model. As in (Zheng et al., 2024), we also report the average sequence entropy as a diversity metric. Specifically, we compute the entropy for the number of tokens for each sequence before it is decoded by the tokenizer and then report the mean entropy value of 5,000 generated sequences.
@@ D.2 ImageNet
In this experiment, we reuse the pre-trained MaskGiT model for all samplers (Chang et al., 2022). The MaskGiT architecture is a Transformer model (Vaswani, 2017) that consists of 24 layers, 8 attention heads, 768 embedding dimension, and 3,072 hidden dimension. Similar to MDLM, this model implements the carry-over unmasked tokens property in the parameterization of $\mathbf{x}_{\theta}$. The full model architecture and training setup details are available in Chang et al. (2022).
In Table 2, the MaskGiT sampler refers to the heuristic confidence-based decoding described in Chang et al. (2022). The MDLM sampler refers to using the outputs of the pre-trained MaskGiT denoising model, then applying the zero-mask parameterization and plugging $\mathbf{x}_{\theta}$ into the posterior from (2). The ReMDM sampler refers to using the same $\mathbf{x}_{\theta}$ as that used with the MDLM sampler, but with the posterior from (8). For ReMDM, we explore the max-capped (with $\eta_{cap}\in\{0.01,0.02,0.05\}$), rescaled (with $\eta\in\{0.01,0.02,0.05\}$), and confidence-based schedules described in Section 4.1. We do not combine these strategies, but rather test each separately. For all three samplers, we perform a sweep over softmax temperature scaling using a scale parameter $\tau\in\{0.6,0.8,1.0\}$. For a vector $\mathbf{y}$, with $\mathbf{y}_{i}$ denoting its $i^{\text{th}}$ component, softmax temperature scaling is implemented as follows:
@@
We use the code from Dhariwal & Nichol (2021) to compute FID and IS metrics.
@@ D.3 QM9
For the guidance experiments on QM9, we follow the setup used in Schiff et al. (2024). The dataset was tokenized using a regular expression-based tokenizer (Schwaller et al., 2019) with padded max sequence lengths of $L=32.$ Including special tokens, the tokenizer vocabulary size is $|V|=40.$ The AR, MDLM, and UDLM models used for discrete CBG and CFG were taken from the implementation provide in Schiff et al. (2024). These models are based on a Transformer architecture (causally masked for AR) known as DiT (Peebles & Xie, 2023). See Schiff et al. (2024) for the full model and experimental setup details. Note that Schiff et al. (2024) provide a first-order approximation for D-CBG that avoids the required $\mathcal{O}(|V|\cdot L\cdot T)$ calls to the classifier $p_{\phi}$, which comes from needing to evaluate every possible token replacement from a vocabulary of size $|V|$ at every position in a sequence of length $L$ for each step of the sampling process of duration $T$. However, given the relatively shorter sequences and smaller vocabulary size used in the QM9 experiments and that Schiff et al. (2024) found generally higher quality results without the first order approximation, we omit the first-order approximation from our experiments and instead use the full, un-approximated D-CBG implementation.
In the main text, we report results for experiments performed to maximize the ring count value of molecules. In Appendix E.4, we also present results for maximizing the property of drug-likeness (QED), which was also explored in Schiff et al. (2024). Guidance training and inference were performed on the molecular property of interest (ring count or QED) using a binarized label, where molecules with property value $\geq$ the $90^{\text{th}}$ percentile in the dataset were labeled $y=1.$ The QED and ring count properties were extracted for each molecule in the QM9 dataset using the RDKit library (Landrum et al., 2013).

### v2

@@ 6 Discussion and Related Work
Denoising Diffusion Implicit Models  An analogy can be drawn between our work and DDIM [48], where ReMDM is to MDLM for discrete signals as DDIM is to DDPM [18] for continuous data. Both our work and DDIM propose alternative forward and backward processes that maintain marginals while enabling more flexible generation. Of note, Song et al. [48] also present a way of adapting processes to discrete domains; however, they focus on uniform categorical as the limiting distribution. In Appendix B, we demonstrate how this can be adapted for absorbing state diffusion and derive an equivalence to our method.
@@
Discrete Predictor-Corrector Samplers  Continuous time Markov chain (CTMC) theory provides a framework for samplers that correct errors in the reverse process, extending the original predictor-corrector formulation for continuous data [50] to discrete domains. In addition to the FB and DFM correctors discussed above, other correctors include: the Stein operator [62], and DPC [22] which improves the sample quality of MaskGiT by corrector sampling. Unlike the plug-and-play methods of FB, DFM, and ReMDM, the Stein operator and DPC correctors require additional training.
@@ 7 Conclusion
In this work, we have presented a novel family of absorbing state discrete diffusion samplers. Our method leverages the strong language modeling performance of this class of models by enabling the use of pretrained weights with the added benefit of more flexible sampling strategies that allow for remasking of predicted tokens. We demonstrate empirically that this leads to improved sample quality for both unconditional and conditional generation and leads to better controlled generation and downstream performance. Our approach also unlocks an important inference-time compute scaling axis that is more limited for existing masked diffusion models.
@@ A.4 Deriving the ReMDM NELBO
The variational objective or negative evidence lower bound (NELBO) for diffusion models has the following form [47]:
@@
Starting from (A.4), we note that, in practice, we set $p_{\theta}(\mathbf{z}_{1})=q(\mathbf{z}_{1}\mid x)=q(\mathbf{z}_{1})=\bm{\pi}$ which ensures $\mathcal{L}_{prior}=0$. Additionally, the reconstruction loss $\mathcal{L}_{recon}$ is equivalent in both (A.4) and (3.2). We therefore turn our attention to the ReMDM diffusion loss term
@@
Additionally, we note that in Sahoo et al. [38], MDLM is parameterized using the denoising $\mathbf{x}_{\theta}$ that is restricted in two ways. First the denoising model places zero probability mass on the [MASK] token, i.e., $\mathbf{x}_{\theta}^{\top}\bm{m}=0.$ Second the model ‘carries-over’ unmasked tokens, i.e., $\mathbf{x}_{\theta}^{\top}\mathbf{z}_{t}=1,$ if $\mathbf{z}_{t}\neq\bm{m}.$ Below we assume that our denoising network follows a similar parameterization.
@@
Then rewriting the summation in (21) as an expectation with $t$ sampled uniformly from $\{\frac{1}{T},\ldots,1\}$ and plugging in (24), we recover the diffusion loss term in (3.2).
@@
The forward-backward corrector sampler [5] is derived using continuous time Markov chain theory. In particular, the rate matrix for the corrector sampler step $\mathbf{R}_{corrector}$ is the sum of the forward rate matrix $\mathbf{R}_{t}$ and the backward rate matrix $\tilde{\mathbf{R}}_{t}$, i.e. $\mathbf{R}_{corrector}=\mathbf{R}_{t}+\tilde{\mathbf{R}_{t}}$.
@@
From Sahoo et al. [38], we know that the forward step of MDLM $q(\mathbf{z}_{t}\mid\mathbf{z}_{s})=\mathrm{Cat}(\mathbf{z}_{t};\frac{\alpha_{t}}{\alpha_{s}}\mathbf{z}_{s}+(1-\frac{\alpha_{t}}{\alpha_{s}})\bm{m})$. Let $\mathcal{V}$ denote the vocabulary set.
@@
The DFM corrector sampler [13] is derived from a generating velocity $u_{t}^{corr}$ using the following transformation equation.
@@
Note that our formulation is slightly different from the original presentation in Gat et al. [13], since in our notation as $t$ goes from 1 to 0, we move from noise to the target distribution, whereas in the flow matching literature this direction is reversed. Plugging (46) and (47) into (45), we derive the following form for $u_{t}^{corr}$:
@@
A.8 Proof of Proposition 4.4
Proof.
Recall that the ReMDM posterior is defined as:
$q_{\sigma}(\mathbf{z}_{s}\mid\mathbf{z}_{t},\mathbf{x})=\begin{cases}\mathrm{Cat}(\mathbf{z}_{s};(1-\sigma_{t})\mathbf{x}+\sigma_{t}\bm{m}),\quad\mathbf{z}_{t}\neq\bm{m}\\ \mathrm{Cat}\left(\mathbf{z}_{s};\frac{\alpha_{s}-(1-\sigma_{t})\alpha_{t}}{1-\alpha_{t}}\mathbf{x}+\frac{1-\alpha_{s}-\sigma_{t}\alpha_{t}}{1-\alpha_{t}}\bm{m}\right),\quad\mathbf{z}_{t}=\bm{m}.\end{cases}$ (51)
When $\alpha_{s}=\alpha_{t}=\alpha$, the posterior simplifies to:
$q_{\sigma}(\mathbf{z}_{s}\mid\mathbf{z}_{t},\mathbf{x})=\begin{cases}\mathrm{Cat}(\mathbf{z}_{s};(1-\sigma_{t})\mathbf{x}+\sigma_{t}\bm{m}),\quad\mathbf{z}_{t}\neq\bm{m}\\ \mathrm{Cat}\left(\mathbf{z}_{s};\frac{\sigma_{t}\alpha}{1-\alpha}\mathbf{x}+\frac{1-(1+\sigma_{t})\alpha}{1-\alpha}\bm{m}\right),\quad\mathbf{z}_{t}=\bm{m}.\end{cases}$ (52)
For $0\leq\sigma_{t}\leq\min\left\{1,\frac{1-\alpha}{\alpha}\right\}$, equation equation 52 defines a valid probability distribution.
In contrast, the DFM sampler applies the following transformation:
$\mathbf{z}_{s_{c}}\sim\delta_{\mathbf{z}_{t}}(\cdot)+u_{t}^{\text{corr}}(\cdot,\mathbf{z}_{t})\Delta t$ (53)
where the corrector velocity $u_{t}^{\text{corr}}$ is a weighted combination of the forward velocity $\hat{u_{t}}$ and backward velocity $\check{u_{t}}$:
$u_{t}^{\text{corr}}(\cdot,\mathbf{z}_{t})=(1+\beta_{t})\hat{u}_{t}(\cdot,\mathbf{z}_{t})-\beta_{t}\check{u}_{t}(\cdot,\mathbf{z}_{t})$ (54)
with
$\hat{u}_{t}(\cdot,\mathbf{z}_{t})$ $=-\frac{\alpha_{t}^{\prime}}{1-\alpha_{t}}\left[p_{0|t}(\cdot\mid\mathbf{z}_{t})-\delta_{\mathbf{z}_{t}}(\cdot)\right],$ (55)
$\check{u}_{t}(\cdot,\mathbf{z}_{t})$ $=-\frac{\alpha_{t}^{\prime}}{\alpha_{t}}\left[\delta_{\mathbf{z}_{t}}(\cdot)-p_{1|t}(\cdot\mid\mathbf{z}_{t})\right].$ (56)
If there exists a time interval $[t,t-\Delta t]$ with $\Delta t>0$ during which $\alpha_{t}$ remains constant, then $\alpha_{t}^{\prime}=0$ throughout this interval. Consequently, both $\hat{u}_{t}$ and $\check{u}_{t}$ vanish, leading to $u_{t}^{\text{corr}}=0$ and hence $\mathbf{z}_{s_{c}}\sim\delta_{\mathbf{z}_{t}}(\cdot)$. In other words, the DFM corrector becomes a degenerate sampler that merely copies the existing token without making any changes.
In conclusion, Proposition 4.3 implies that the DFM corrector is a special case of the more general ReMDM sampler. The above analysis further establishes that this inclusion is strict. Thus, ReMDM strictly generalizes the DFM corrector.
@@ Appendix B Comparison to DDIM Non-Markovian Processes
In DDIM [48], the authors present non-Markovian forward processes for both continuous and discrete signals. For discrete data, although in DDIM the proposed method assumes a uniform categorical as the limiting distribution, here we demonstrate how one can adapt the proposed method in DDIM for absorbing state diffusion and derive an equivalence between ReMDM and this new DDIM process under a reparameterization of $\sigma_{t}$. For clarity, we use $\sigma_{t}^{DDIM}$ to denote the parameter from DDIM and $\sigma_{t}^{ReMDM}$ to denote the parameter used in our work.
@@ Appendix C ReMDM Sampler Algorithms
Appendix D Additional Experimental Details
D.1 OpenWebText
In this experiment, we reuse the pretrained AR, SEDD, and MDLM checkpoints released by [38] where the diffusion models are trained using a log-linear schedule, i.e., $\alpha_{t}=1-t$. AR, SEDD, and MDLM share the same architecture: a Transformer-based model [53] that augments the diffusion transformer [32] with rotary embeddings [52] and consists of 169M parameters. The neural network is comprised of 12 layers, 12 attention heads, and 768 hidden dimensions. Please see Sahoo et al. [38] for the full model architecture and training details.
@@ D.1 OpenWebText
As in Sahoo et al. [38], we use gpt-2 tokenizer for our experiments. We use the same train-validation split as in Sahoo et al. [38] (where the last 100k documents of OWT were designated as the validation set) and randomly select 5,000 samples from the validation set to serve as the ‘reference’ for MAUVE score computation. For the discrete flow matching (DFM) corrector sampler, we follow Gat et al. [13] and set the corrector schedule $\beta(t)=At^{0.25}(1-t)^{0.25}$ where $A=10$ (see Appendix A.7 for the definition of corrector schedule). We empirically find that nucleus sampling performs better than the temperature technique proposed in Gat et al. [13]. We also use non-fixed-width time steps as in Gat et al. [13].
For evaluation metrics, we report MAUVE scores, generative perplexity, and entropy. For MAUVE, we generate 5,000 samples for each model/sampler. We use the gpt-2 tokenizer, GPT-2 Large [35] as the embedding model, and the MAUVE scaling hyperparameter is set to 5. For generative perplexity, we use GPT-2 Large as the external model. As in [63], we also report the average sequence entropy as a diversity metric. Specifically, we compute the entropy for the number of tokens for each sequence before it is decoded by the tokenizer and then report the mean entropy value of 5,000 generated sequences.
@@ D.2 ImageNet
In this experiment, we reuse the pre-trained MaskGiT model for all samplers [7]. The MaskGiT architecture is a Transformer model [53] that consists of 24 layers, 8 attention heads, 768 embedding dimensions, and 3,072 hidden dimensions. Similarly to MDLM, this model implements the carry-over unmasked tokens property in the parameterization of $\mathbf{x}_{\theta}$. The full model architecture and training setup details are available in Chang et al. [7].
In Table 2, the MaskGiT sampler refers to the heuristic confidence-based decoding described in Chang et al. [7]. The MDLM sampler refers to using the outputs of the pre-trained MaskGiT denoising model, then applying the zero-mask parameterization and plugging $\mathbf{x}_{\theta}$ into the posterior from (2). The ReMDM sampler refers to using the same $\mathbf{x}_{\theta}$ as that used with the MDLM sampler, but with the posterior from (51). For ReMDM, we explore the max-capped (with $\eta_{cap}\in\{0.01,0.02,0.05\}$), rescaled (with $\eta\in\{0.01,0.02,0.05\}$), and confidence-based schedules described in Section 4.1. We do not combine these strategies, but rather test each separately. For all three samplers, we perform a sweep over softmax temperature scaling using a scale parameter $\tau\in\{0.6,0.8,1.0\}$. For a vector $\mathbf{y}$, with $\mathbf{y}_{i}$ denoting its $i^{\text{th}}$ component, softmax temperature scaling is implemented as follows:
@@
We use the code from Dhariwal & Nichol [9] to compute FID and IS metrics.
@@ D.3 QM9
For the guidance experiments on QM9, we follow the setup used in Schiff et al. [43]. The dataset was tokenized using a regular expression-based tokenizer [44] with padded max sequence lengths of $L=32.$ Including special tokens, the tokenizer vocabulary size is $|V|=40.$ The AR, MDLM, and UDLM models used for discrete CBG and CFG were taken from the implementation provide in Schiff et al. [43]. These models are based on a Transformer architecture (causally masked for AR) known as DiT [32]. See Schiff et al. [43] for the full model and experimental setup details. Note that Schiff et al. [43] provide a first-order approximation for D-CBG that avoids the required $\mathcal{O}(|V|\cdot L\cdot T)$ calls to the classifier $p_{\phi}$, which comes from needing to evaluate every possible token replacement from a vocabulary of size $|V|$ at every position in a sequence of length $L$ for each step of the sampling process of duration $T$. However, given the relatively shorter sequences and smaller vocabulary size used in the QM9 experiments and that Schiff et al. [43] found generally higher quality results without the first order approximation, we omit the first-order approximation from our experiments and instead use the full, un-approximated D-CBG implementation.
In the main text, we report results for experiments performed to maximize the ring count value of molecules. In Appendix E.4, we also present results for maximizing the property of drug-likeness (QED), which was also explored in Schiff et al. [43]. Guidance training and inference were performed on the molecular property of interest (ring count or QED) using a binarized label, where molecules with property value $\geq$ the $90^{\text{th}}$ percentile in the dataset were labeled $y=1.$ The QED and ring count properties were extracted for each molecule in the QM9 dataset using the RDKit library [21].
@@
D.4 LLaDA
We reuse the open-source LLaDA 8B Instruct model from [28] in this experiment. Since LLaDA utilizes MaskGiT-style samplers, we modify ReMDM-loop to similar samplers with confidence heuristics. For a maximum mask length equal to 32, we first use the LLaDA sampler to generate 28 tokens. Then we step into the ReMDM-loop phase, where we remask a fixed number (1 for Countdown and 4 for TruthfulQA) of tokens with the smallest decoding confidence and unmask the same number of tokens each time step. The ReMDM-loop phase lasts 32 steps, and after the loop phase, we use the LLaDA sampler to fill in the remaining gaps. For DFM, since it does not support looping (see Proposition 4.4), we remove the loop phase and let it remask one token and unmask two tokens each time step instead. Different from LLaDA, which utilizes greedy sampling, we enable random sampling to compute error bars.
We use the widely used evaluation harness [12] as the evaluation framework, and for both Countdown and TruthfulQA, we use a maximum length of 32 tokens and no semi-autoregressive sampling introduced in [28], that is, the block size equals 32 due to relatively short answer lengths. For Countdown, we use 4-shot evaluation, and for TruthfulQA, we use 6-shot evaluation (see Appendix F.3 for the chat templates we use). For Countdown, we evaluate on 256 synthetically generated countdown questions. We repeat the sampling process for each question 10 times and calculate pass@1 as $\frac{c}{10}$ where $c$ is the number of correct trials. After this, we report the average pass@1 for the 256 questions. For TruthfulQA, we let the model answer 817 questions on factual knowledge and calculate the ROUGE scores between the model and correct answers, as well as the model and plausibly wrong answers. The difference between the right and wrong ROUGE scores is reported. Due to the low computational resource demand for the two tasks (one run takes around 30 minutes using one Nvidia A100 GPU), for both tasks, we repeat the experiments 20 times with 20 different random seeds (0-19) and report the mean values and 95% confidence intervals.